OpenStudy (anonymous):

Evaluate. 8 – 2 • 3 + 7

OpenStudy (anonymous):

is it 25 

OpenStudy (anonymous):

PEMDAS=parenthesis, exponents, multiplication, division, addition then subtraction this is the order of evaluating the operations. start with the multiplication first, then add, then subtract

OpenStudy (anonymous):

\[=8-2\times3+7\] \[=8-6+7\] \[=?\] solve it @shelbydancer

OpenStudy (anonymous):

9 @MARC_

OpenStudy (anonymous):

good :)